OpenStudy (anonymous):

Evaluate the following expression using the values given: Find -a2 - 3b3 + c2 + 2b3 - c2 if a = 3, b = 2, and c = -3. a.-67 b.-17 c.-31 d.1 PLEASE HELP!!!

OpenStudy (anonymous):

Will give medal

OpenStudy (anonymous):

Here you go Let's simplify step-by-step. −a2−3b3+c2+2b3−c2 =−a2+−3b3+c2+2b3+−c2 Combine Like Terms: =−a2+−3b3+c2+2b3+−c2 =(−3b3+2b3)+(−a2)+(c2+−c2) =−b3+−a2 Answer: =−b3−a2

OpenStudy (anonymous):

they dont have that as a choice

OpenStudy (anonymous):

You have to plug in the numbers

OpenStudy (anonymous):

I think it is A but I am not sure

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

you were wrong